If f(1) 4 and f' (1)= 2 .find the derivative of log f(e power x) at x= 0......also how to find f(x) from the question given

Grade:12

1 Answers

Saswat Pani
6 years ago
F(1)= 4,F`(1)=2A.T.Q=d/dt(log f(e^x)) at x=0=1/f(e^x)×f`(e^x)Putting x=0=1/f(1)×f`(1)=1/4×2=0.5